OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3

OpenStudy (angela210793):

When you are given a point and the slope u can find the equation of the line by using this formula: y-y0=m*(x-x0) where (x0;y0) are the coordinates of the point given and m is the slope so u will have y-(-8)=3(x-8) y+8=3*x-3*8 y+8=3x-24 y=3x-24-8 y=3x-32